OpenStudy (anonymous):

is y=3x^2-4x+5 y=(x-2)^2-3 y=-2x^2-9 and y=-4(x-2)(x+3) either a general form, vertex form, factor form or none of these? How can I tell?

OpenStudy (anonymous):

General form is y=ax^2+bx+c vertex form is y=a(x-h)^2+k

OpenStudy (anonymous):

So the first one is a general and the second and last one is vertex

OpenStudy (anonymous):

Also is there a way that I can convert the ones that aren't in general form into general form?

OpenStudy (anonymous):

Last one isnt vertex form

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

y=-2x^2-9 is general form y=-4(x-2)(x+3) is factored form to get them into general form just multiply them out

OpenStudy (anonymous):

So how would I do that exactly?????

OpenStudy (anonymous):

for y=(x-2)^2-3 y=(x-2)(x-2)-3 foil it y=x^2-4x+4-3 combine like terms y=x^2-4x-1

OpenStudy (anonymous):

hmm okay